b. What is the IRR of the project? C1 -$6,750 +$4,500 $18,000 CoPlease solve without using excel:
Answer #1
0 1 2
-6750 4500 18000
IRR= 100%

Only method of calculating IRR is by trial and error
Solve following equation for NPV=0

NPV= C0/(1+r)^0+C1/(1+r)^1+C2/(1+r)^2
Take r= 50% =-6750/(1+50%)^0+4500/(1+50%)^1+18000/(1+50%)^2
4250
Take r= 125% =-6750/(1+125%)^0+4500/(1+125%)^1+18000/(1+125%)^2
-1194.44
Take r= 100% =-6750/(1+100%)^0+4500/(1+100%)^1+18000/(1+100%)^2
0.00

The choice of guess was influenced by the fact the we knew what the expected answer should be based on the excel calculator. However, the general method is the same. When NPV comes to +ve, increased the discount rate estimate; when the NPV comes -ve, decrease the discount rate estimate and so on